The velocity of the Rio GRand River is 3mph. Moving with the current a canoe can travel 10miles in the same amount of time it would take to go 4 miles moving against the current. Find the canoe's rate in still water.
Answer by richwmiller(17219)

g=3
r*t=d
t=d/r
(r+g)*t=10
(r-g)*t=4
t=10/(r+3)
t=4/(r-3)
10/(r+3)=4/(r-3)
r=7
check
7+3*t=10
7-3*t=4
ok